OpenStudy (anonymous):

PLEASE HELP :3 Samantha is going on vacation for the summer and is trying to choose between two different plans. The first plan costs $450 for 3 days at a hotel and 2 days at an amusement park. The second plan offers 5 days at the same hotel and 3 days at the amusement park for $700. The cost of 1 day at the hotel and the amusement park is the same under both plans. How much does a 1 day trip to the amusement park cost? $50 $75 $150 $200

OpenStudy (anonymous):

Solve 7x − 2y = −3 14x + y = 14 (4, 5/7) (4, 7/5) (5/7, 4) (7/5, 4)

OpenStudy (anonymous):

@mhmdrz91

OpenStudy (anonymous):

Solve x + 3y = 9 3x − 3y = −13 (−1, 10/3) (1, -10/3) (−1, 3/10) (1, -3/10)

OpenStudy (anonymous):

7x - 2y = -3 => 7x = 2y-3 14x + y = 14 => 7x +y/2 = 7 => 7x = 7-y/2 now 2y-3 = 7-y/2 2y+y/2 = 10 -> 5y/2 = 10 -> 5y = 20 -> y = 4 7x-8=-3 -> 7x = 5 -> x = 5/7 answer is (5/7,4)

OpenStudy (anonymous):

for x+3y=9 and 3x - 3y = -13 use same as last solution ;-)

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

What do u mean last solution?

OpenStudy (anonymous):

7x - 2y = -3 14x + y = 14 solution for this :D

OpenStudy (anonymous):

Oh.. im confused.. I do the same thing for the last one that u did for the 7x one?

OpenStudy (anonymous):

x + 3y = 9 => 3x = 27 - 9y 3x − 3y = −13 => 3x = 3y -13 27-9y = 3y -13 -> 12y = 40 -> y = 40/12 = 10/3 x + 3*(10/3) = 9 -> x = 9 - 10 = -1 (-1 , 10/3) is answer